OpenStudy (chillhill):

On a given day, machine A has a 10% chance of malfunctioning and machine B has a 7% chance of the same. Given that at least one of the machines malfunctioned today, what is the chance that machine B malfunctioned?

OpenStudy (anonymous):

They are asking for \(\Pr(B|A\cup B)\). They are given you \(\Pr(A)\) and \(\Pr(B)\).

OpenStudy (anonymous):

You're going to want to use: \[ \Pr(B|A\cup B) = \frac{\Pr(B\cap (A\cup B))}{\Pr(A\cup B)} \]

OpenStudy (anonymous):

Now, since \(B\subseteq A\cup B\), we can say \(B\cap (A\cup B)=B\). Thus we can say: \[ \Pr(B|A\cup B) = \frac{\Pr(B)}{\Pr(A\cup B)} \]

OpenStudy (anonymous):

Next we use our formula for disjoint probability: \[ \Pr(A\cup B) = \Pr(A)+\Pr(B)-\Pr(A\cap B) \]Since we have no reason not to think that the machines fail independently, we can use the formula for independent random events: \[ \Pr(A\cap B) = \Pr(A)\Pr(B) \]Therefore: \[ \Pr(A\cup B) = \Pr(A)+\Pr(B)-\Pr(A)\Pr(B) \]

OpenStudy (anonymous):

Finally, \[ \Pr(B|A\cup B) = \frac{\Pr(B)}{\Pr(A)+\Pr(B)-\Pr(A)\Pr(B)} \]Since they gave us \(\Pr(A)\) and \(Pr(B)\), we have our answer.